OpenStudy (anonymous):

d/dx ln(x2+x) find the indicated derivative and simply. The ln(x2+x) is under a radical

OpenStudy (anonymous):

\[d/dx \sqrt{\ln(x^{2}+x)}\]=(ln(x^{2}+x)^(1/2) Then you can use the rule of powers to find the derivative. The rule of powers says d/dx (u^n)=n(u^(n+1)). The derivative of ln v=1/v. By combining these two rules you can find the derivative of your function

OpenStudy (anonymous):

Thank you